OpenStudy (anonymous):

Little confused on this? Just working on parametric equations, I have x= sin theta/2 & y = cos theta/2... and I have to create a plot table between -pie and pie. I don't know if Im supposed to use the radians as the parameter for the plot table or just degrees? i'm confused...

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

always use radians, they are numbers

OpenStudy (misty1212):

if it is \[x=\sin(\frac{\theta}{2}), y=\cos(\frac{\theta}{2})\] you should get a circle

OpenStudy (misty1212):

actually because of the \(\frac{\theta}{2}\) you get the upper half of a circle

OpenStudy (anonymous):

right, I see how it's a circle, i'm just a little confused on the radians, because I'm not understanding what values are supposed to correspond on the table?... I'm sorry if this is confusing, because I'm barely refreshing my trig values and stuff, and I'm not understanding just what values do I use as the parameter...

OpenStudy (anonymous):

so would my value -pie correspond to -1? say for like the first value on my table?... if I was just plugging in the value for sin 0/2?

zepdrix (zepdrix):

I guess the start of your interval, \(\rm \theta=-\pi\) would correspond to \(\rm (x,y)=(-1,0)\) yes?

OpenStudy (misty1212):

here is a nice picture this is what you should get http://www.wolframalpha.com/input/?i=x%3Dsin%28t%2F2%29%2Cy%3Dcos%28t%2F2%29%2C+t+from+-pi+to+pi

OpenStudy (anonymous):

ok. great. So If i was plotting the points for the table and I started at -pi, and then -3pi/2, -pi/2, etc. up to pi. Would I just plug in the corresponding degree into the function to be able to graph it by hand.. for example, x= sin theta/2 I would plug in for first value sin 1/2 to get my x coordinate?

OpenStudy (misty1212):

at \(-\pi\) you get \[(\sin(-\frac{\pi}{2}),\cos(-\frac{\pi}{2}))=(0,-1)\]

zepdrix (zepdrix):

Ahh woops :) I just assumed sine went with the y coordinate haha

OpenStudy (misty1212):

yeah it usually does, doesn't it?

zepdrix (zepdrix):

Oh maybe I didn't... sin(-pi/2) = -1

zepdrix (zepdrix):

Ya I think misty has those backwards :d

OpenStudy (anonymous):

ok. great. I see what your saying, so If I was to plug in the value for -3pi/2 do I just convert that into degrees and make it sin -3pi/2 /2 ?

OpenStudy (misty1212):

lol totally backwards

OpenStudy (misty1212):

i did the cosine first and the sine second, same mistake doe!

zepdrix (zepdrix):

degrees? ew ew ew :( \[\large\rm \frac{3\pi/2}{2}=\frac{3\pi}{4}\]

OpenStudy (anonymous):

oh. ok. I understand now. great. thank you, Yes I was not understanding the x y points. thank you